OpenStudy (anonymous):

[1/1+6] + [3/x+4] = -2 / x^2+10x+24

OpenStudy (anonymous):

kathy, something is missing, is it [1/1+6x]?

OpenStudy (anonymous):

sorry , 1 / x+6

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

Again find what are the denominator for each of the fractions.

OpenStudy (anonymous):

on the left hand side, you have x+6 and x+4. What do you have on the right side?

OpenStudy (anonymous):

x^2+ 10 + 24 = (x+6)(x+4)

OpenStudy (anonymous):

but how do i put the other together to make the same

OpenStudy (anonymous):

great kathy. So we only have z+6 and z+4 in the denominator. On the right side, we don't have to do anything. For the left side, make each fractions to have (x+6)(x+4) as denominator. Then compare the numerator

OpenStudy (anonymous):

oops...typo..x+6 and x+4

OpenStudy (anonymous):

[1/x+6] + [3/x+4] = -2 / x^2+10x+24 [(x+4)/(x+6)(x+4)] + [3(x+6)/(x+4)(x+6)] = -2 /(x+6)(x+4)

OpenStudy (anonymous):

hmmm i dont get it, we dont have to make the left side denominator the same as the right ?

OpenStudy (anonymous):

oh i see, but what do i do after?

OpenStudy (anonymous):

compare the numerator, because now you already have the denominator to be the same. The idea is just the same when we deal with number. I hope the following illustration with numbers can help you: 1/2 + 1/4 = y/24 12/24 + 6/24 = y/24 -> 12 + 6 = y

OpenStudy (anonymous):

oh so now i just add : x+4 +3(x+6) = -2

OpenStudy (anonymous):

/ (x+6)(x+4)

OpenStudy (anonymous):

yup, you are right kathy. well done!

OpenStudy (anonymous):

[x+4 +3(x+6)]/(x+6)(x+4) = -2/(x+6)(x+4) by comparing the numerator, you get x+4 +3(x+6) = -2

OpenStudy (anonymous):

4x+22=-2/ (x+6)(x+4)

OpenStudy (anonymous):

no need the / (x+6)(x+4)

OpenStudy (anonymous):

so it cancels out ?

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

but what happens to the denominator

OpenStudy (anonymous):

I wont say cancel out, I would rather say for the left and right to be equal and since the denominators are same, then the numerator must be the same

OpenStudy (anonymous):

[x+4 +3(x+6)]/(x+6)(x+4) = -2/(x+6)(x+4) by comparing the numerator, you get x+4 +3(x+6) = -2

OpenStudy (anonymous):

which equals to : 4x + 22 = -2 4x = 20 x=5

OpenStudy (anonymous):

4x + 22 = -2 4x = -24 x = -6

OpenStudy (anonymous):

my mistake. so after that what would happen to the denominator so we can find x

OpenStudy (anonymous):

well, the original equation is [1/x+6] + [3/x+4] = -2 / x^2+10x+24 from this we know that our solution x cannot be -6 or -4. So as a conclusion, i will say there is no solution for this problem, even though we solve to get x = -6

OpenStudy (anonymous):

Does it make sense to you?

OpenStudy (anonymous):

true, so it has no answer ?

OpenStudy (anonymous):

ya, there is no answer for this problem

OpenStudy (anonymous):

1/2t + 6/t =1 1/2t + 12/2t =13/2t 13/2t = 1 divide both by 2 to leave x alone & its 13/2

OpenStudy (anonymous):

just doube checking this other problem

OpenStudy (anonymous):

or is it 7/2

OpenStudy (anonymous):

are you still there?